INAF LBC+MODS queue observing, 2012 February 22-23

Observer: K. Boutsia, M. Fumana
Telescope Operator: G. Bechetti
Support Astronomer: O. Kuhn
Instrument: LBC+MODS

Summary

Some thin clouds in the west during sunset. Probably clouds passing by during the whole night. Seeing has been very poor.
We started with LBC but the seeing was really bad (2.0"-3.0") so we did not complete the NSI observations.
After that we switched to MODS. There have been problems with the swing arms and then we had problems collimating (due to bad seeing?). 1hr16min downtime.
With MODS we obtained data for Nonino (0647), Fiore (mask2) and Donnarumma (A1703).
For GCS seeing measurements during each program see figures at the end of the log.
Calibration data for MODS images can be found in Repository for 2012022UT, 20120223UT and 20120220UT.

Data Summary

Project Objects Filter ExpTime Notes
Valsecchi 1999AO10 g-sloan / r-sloan 30min seeing was bad and we did not go on with the second pointing
Storm IC10VI V-bess / I-bess 5min seeing was >3.0", measured on the images, so we aborted
Nonino macsj0647 dual grating 1hr seeing was >1.5"
Fiore Cosmos2 dual grating 3hr seeing was ~1.7" but this is comparable to yesterdays data - we have 6.5hr total - completed (?)
Donnarumma A1703m1 red grating 1hr15min seeing was around ~1.2" for most of the time going up to 1.4" at the end
STD_MODS Feige34 dual/red grating good for all MODS programs

Details

01:53 Slew to 20120223F
  • dofpia,/First,/X2
15199      0      0      0      0      0      0   1.27   5.99     red
 18170      0      0      0      0      0      0   1.49   7.16    blue

-173    -24     79    -44     59    255    100   2.16   0.20     red
 453   -213    -45   -114    207    187     96   2.38   0.26    blue
  • lbcrangebal
  • it is still bright outside, so we run dofpia again and wait for the dimm
-2166    -33     99    161   -185    270    100   2.04   0.88     red
-1197   -113      6     92    -25    232     97   2.16   0.50    blue

-678   -104     70    113   -104    150     92   1.94   0.30     red
-519    -56    105    -67    -50    206     83   2.27   0.26    blue

  32    -46    -68   -291    109     68     72   1.88   0.17     red
 130    -80     83      3     17    101     52   1.99   0.12    blue
  • dimm 2.1"

02:15 Take an image of the first offset position to see what the seeing is in the field.
  • seeing on images 1.6". Try with observation

02:20 Start with 201202122E

03:04 dofpia on 201202122W
 -8554      0      0      0      0      0      0   1.64   3.37     red
-8164      0      0      0      0      0      0   1.68   3.22    blue

 40    -44   -193    137     42    136     50   1.88   0.15     red
-196   -110    -18    287   -470    301     74   2.15   0.33    blue
  • seeing measured on the image is 3.0" - seeing from dofpia 2.0"
 -753    185    -43     26    -14     92     33   1.66   0.32     red
 -734   -151    -21     26     36    101     44   1.81   0.31    blue

-144      3     85   -101    174    165     51   1.97   0.17     red
-260     -6   -112    124     -5    182     54   1.95   0.18    blue
  • dimm seeing 2.45"

03:25 slew to IC10VI - airmass 2.12
  • dofpia,/X2
-1228   -153     -5   -405    376      0      0   2.58   0.55     red
-2441   -158    -33   -647    272      0      0   2.86   1.02    blue

69    -35    118    203    -58    334     88   2.41   0.25     red
 -91   -155    -73    182   -132      0      0   2.50   0.14    blue

03:31 IC10VI
  • seeing measured on the images >3.0"
  • abort OB

03:40 SWITCH TO MODS

04:36 We are on a star collimating but it does not look good.
  • seeing on the guider 2.6"

05:06 acqMODS macsj0647_mods.acq mods1r 0003-0005
  • seeing on guider 2.0"
  • offsetpointing -0.1222 -0.118 -0.880 detxy rel
  • dimm seeing 2.2"
  • seeing measured on the acquisition image was 1.6"

05:21 execMODS macsj0647_mods.obs mods1b 0003 - 0006 / mods1r 0006-0009

06:33 acqMODS mods.57.FIORE2.acq mods1r 0010-0012
  • offsetpointing -0.2532 0.112 -0.868 detxy rel
  • the ref sources on the upper part of the mask look a bit off on the x-axis and we repeat acquisition

06:50 acqMODS mods.57.FIORE2.acq mods1r 0013-0015
  • this time it looks better
  dX:   0.132 arcsec
   dY:  -1.106 arcsec
  dPA: -0.1733 degrees

07:00 execMODS mods.57.FIORE2.obs mods1b 0007-0012 / mods1r 0016-0021
  • maybe also some clouds passing by
  • on the guider we see that a lot of times it looses the guide star (there is nothing there - fhwm 0.0)

10:14 Abort - we finished 3hr with this mask

10:14 acqMODS mods.56.FIORE1.acq mods1r 0022-0024
  dX:  -0.154 arcsec
   dY:  -0.714 arcsec
  dPA: -0.1422 degrees

10:26 execMODS mods.56.FIORE1.obs - ABORT no images saved
  • problems with the IMCS on the red side... it failed twice to lock-on - we lost 15min
  • for long times we have no guide star
  • airmass is reaching the limit - especially because of the delays with the laser failures

10:48 abort FIORE1

10:48 acqMODS Feige34_std.acq mods1r 0025-0029

11:04 execMODS Feige34_std.obs mods1b 0013 / mods1r 0030
  • this is with dual_grating for FIORE

11:10 execMODS Feige34_stdRED.obs mods1r 0031
  • this is with red_grating for Donnarumma

11:16 acqMODS mods.201151.newmi1e.acq mods1r 0032-0034
   dX:  -0.009 arcsec
   dY:  -1.054 arcsec
  dPA: -0.1089 degrees

11:29 execMODS mods.201151.newmi1e.obs mods1r 0035-0036
  • dimm seeing 1.6"

12:49 Abort script after the 2nd exposure because we are already 20min within the twilight.

12:50 Closing dome and going to the zenith for flats with the calibration unit.
  • execMODS -u grmosflats.56.FIORE1.cal mods1b 0014-0018 / mods1r 0037-0041 this script was aborted because the pc got stuck after finishing the first lamp
  • execMODS -u grmosflats.56.FIORE1new.cal mods1b 0019-0023 / mods1r 0042-0046
  • execMODS -u grmoslamps.56.FIORE1.cal mods1b 0024-0026 / mods1r 0047-0049
  • execMODS -u grmosflats.56.FIORE1.cal mods1b 0027-0031 / mods1r 0050-0054 doing the lamp that was not obtained before - red is saturated because no ND1.5 in

02:05 End of calibrations - put MODS to sleep
  • execMODS ~MODSeng/modsScripts/Support/modsSleep.pro

END OF NIGHT

  • Seeing reported by GCS while we were obtaining Nonino 0647:
    nonino.jpg

  • Seeing reported by GCS while we were obtaining FIORE2:
    fiore2.jpg

  • Seeing reported by GCS while we were obtaining Donnarumma:
    dona2.jpg
Topic revision: r6 - 24 Feb 2012, KonstantinaBoutsia
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback